OpenStudy (anonymous):

Factor the following. a^4 − b^4 @thewonderfuladele

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Have you heard about a difference of two squares?

OpenStudy (skullpatrol):

$$\Huge a^2-b^2$$

OpenStudy (anonymous):

uhh, not I can think of... so, it's just A^2-B^2?

OpenStudy (solomonzelman):

@cookiibabii93 do you know what "i" is? or not?

OpenStudy (anonymous):

@cookiibabii93 im going to go now :P other people will help you out surely. Nice meeting you buddy.

OpenStudy (skullpatrol):

$$\Huge a^4=(a^2)^2$$

OpenStudy (solomonzelman):

I have to go off soon. l made my own comp log off at 10:30

OpenStudy (anonymous):

i? you mean like real numbers or something like that?

OpenStudy (anonymous):

Nice meeting you too, @thewonderfuladele , you're fun lol

OpenStudy (skullpatrol):

$$\Huge (a^m)^n=a^{m*n}$$

OpenStudy (skullpatrol):

$$\Huge a^2-b^2=(a+b)(a-b)$$

OpenStudy (anonymous):

So, how would I work this out?

OpenStudy (anonymous):